at what speed does a gun recoil if it weight 44.5 n and fires a 1.0g bullet at a 2500m/s


1
Expert's answer
2022-04-27T13:33:48-0400

The law of conservation of the momentum says

"0=m_bv_b-m_gv_g"

Hence, the speed of the gun after shooting

"v_g=v_b\frac{m_b}{m_g}\\
=2500\:\rm m/s*\frac{1.0*10^{-3}\:kg}{4.54\:kg}=0.55\: m/s"